OpenStudy (anonymous):

This polynomial is a perfect square trinomial: 100x^2-140x+49 True False

OpenStudy (anonymous):

When in doubt, go to the formula: (a-b)^2 = a^2 - 2ab + b^2 With this formula, a would have to be 10x. b would be 7. -2ab = -140x -2(10x)(7) = -140x -140x = -140x This is indeed a perfect square